# Cross-correlation between the curvature perturbations and magnetic fields in pure ultra slow roll inflation

### v1

@@ B.2 Case of $p=4$
$\quad-\,\left[9\,k_{1}^{4}-6\,k_{1}^{2}\,(k_{2}^{2}+k_{3}^{2})-3\,k_{2}^{4}-2\,k_{2}^{2}\,k_{3}^{2}-3\,k_{3}^{4}\right]\,\eta_{\mathrm{e}}^{4}\;\mathrm{Ei}[i\,(k_{1}+k_{2}+k_{3})\,\eta_{\mathrm{e}}]\biggr{)},$ (84d)
${\mathcal{G}}_{5}({\bm{k}}_{1},{\bm{k}}_{2},{\bm{k}}_{3})$ $=\frac{i\,H_{{}_{\mathrm{I}}}}{32\,M_{{}_{\mathrm{Pl}}}\,\sqrt{k_{1}\,k_{2}\,k_{3}^{5}\,\epsilon_{1\mathrm{i}}}}\,\left(\frac{\eta_{\mathrm{i}}}{\eta_{\mathrm{e}}}\right)^{2}\,\left[2\,k_{3}^{2}\,(k_{1}^{2}+k_{2}^{2})+(k_{1}^{2}-k_{2}^{2})^{2}-3\,k_{3}^{4}\right]$
$\quad+(3\,k_{1}^{2}-3\,k_{2}^{2}-k_{3}^{2})\,\eta_{\mathrm{e}}^{2}\;\mathrm{Ei}[i\,(k_{1}+k_{2}+k_{3})\,\eta_{\mathrm{e}}]\bigg{\}}+{\bm{k}}_{2}\leftrightarrow{\bm{k}}_{3},$ (84e)
${\mathcal{G}}_{6}({\bm{k}}_{1},{\bm{k}}_{2},{\bm{k}}_{3})$ $=\frac{i\,H_{{}_{\mathrm{I}}}^{2}\,\epsilon_{1\mathrm{i}}}{16\,M_{{}_{\mathrm{Pl}}}\,\sqrt{k_{1}^{5}\,k_{2}\,k_{3}^{5}\,\epsilon_{1\mathrm{i}}}\,(k_{1}+k_{2}+k_{3})^{5}\,\eta_{\mathrm{i}}^{2}}\,\left[2\,k_{3}^{2}\,(k_{1}^{2}+k_{2}^{2})+(k_{1}^{2}-k_{2}^{2})^{2}-3\,k_{3}^{4}\right]\,\mathrm{e}^{i\,(k_{1}+k_{2}+k_{3})\,\eta_{\mathrm{e}}}$
×$\quad\times\bigg{\{}-2\,i\,\biggl{[}4\,k_{3}^{2}\,(7\,k_{1}^{2}+20\,k_{1}\,k_{2}+7\,k_{2}^{2})+15\,k_{3}\,(k_{1}+k_{2})\,(k_{1}^{2}+3\,k_{1}\,k_{2}+k_{2}^{2})$
$\quad+3\,(k_{1}+k_{2})^{2}\,(k_{1}^{2}+3\,k_{1}\,k_{2}+k_{2}^{2})+20\,k_{3}^{3}\,(k_{1}+k_{2})+4\,k_{3}^{4}\biggr{]}$
$\quad-\biggl{[}2\,k_{3}^{2}\,(19\,k_{1}^{2}+62\,k_{1}\,k_{2}+19\,k_{2}^{2})+6\,k_{3}(k_{1}+k_{2})\,(3\,k_{1}^{2}+11\,k_{1}\,k_{2}+3\,k_{2}^{2})$
$\quad+3\,(k_{1}+k_{2})^{2}\,(k_{1}^{2}+4\,k_{1}\,k_{2}+k_{2}^{2})+28\,k_{3}^{3}\,(k_{1}+k_{2})+5\,k_{3}^{4}\bigg{]}\,(k_{1}+k_{2}+k_{3})\,\eta_{\mathrm{e}}$
$\quad+i\,\biggl{[}2\,k_{3}^{2}\,(5\,k_{1}^{2}+22\,k_{1}\,k_{2}+5\,k_{2}^{2})+3\,k_{3}\,(k_{1}+k_{2})\,(k_{1}^{2}+7\,k_{1}\,k_{2}+k_{2}^{2})$
$\quad+8\,k_{3}^{3}\,(k_{1}+k_{2})+3\,k_{1}\,k_{2}\,(k_{1}+k_{2})^{2}+k_{3}^{4}\bigg{]}\,(k_{1}+k_{2}+k_{3})^{2}\,\eta_{\mathrm{e}}^{2}$
$\quad+\left[3\,k_{1}\,k_{2}\,(k_{1}+k_{2})\,k_{3}+(k_{1}^{2}+9\,k_{1}\,k_{2}+k_{2}^{2})\,k_{3}^{2}+(k_{1}+k_{2})\,k_{3}^{3}\right]\,(k_{1}+k_{2}+k_{3})^{3}\,\eta_{\mathrm{e}}^{3}$
$\quad-i\,k_{1}\,k_{2}\,k_{3}^{2}\,(k_{1}+k_{2}+k_{3})^{4}\,\eta_{\mathrm{e}}^{4}\bigg{\}}+{\bm{k}}_{2}\leftrightarrow{\bm{k}}_{3}.$ (84f)
We should mention that the expressions listed above have been arrived at without any approximations. In Sec. V, we have used these ${\mathcal{G}}_{C}({\bm{k}}_{1},{\bm{k}}_{2},{\bm{k}}_{3})$ to evaluate the contributions ${\mathcal{B}}_{C}{\bm{k}}_{1},{\bm{k}}_{2},{\bm{k}}_{3})$. We should add that, in listing the different contributions ${\mathcal{B}}_{C}({\bm{k}}_{1},{\bm{k}}_{2},{\bm{k}}_{3})$, we have retained only the leading order terms in $\eta_{\mathrm{e}}$.
@@ Appendix C Non-Gaussian contributions to $\mathcal{P}_{{}_{\mathrm{B}}}(k)$
We should mention here that similar analyses in the case of the contributions to the scalar power spectrum $\mathcal{P}_{{}_{\mathrm{S}}}(k)$ due to the scalar non-Gaussianity parameter $f_{{}_{\mathrm{NL}}}({\bm{k}}_{1},{\bm{k}}_{2},{\bm{k}}_{3})$ has been carried out earlier [113, 114, 48, 49, 115]. We shall suitably adopt the method followed in the case of the scalar power spectrum (in this context, see Refs. [49, 115]). Using the defining relation (55), we can express the power spectrum of the magnetic field that is modified due to the non-Gaussianity parameter $b_{{}_{\mathrm{NL}}}({\bm{k}}_{1},{\bm{k}}_{2},{\bm{k}}_{3})$ to be
@@
In arriving at the second equation we have used the fact that $\mathcal{P}_{{}_{\mathrm{B}}}(k)=\mathcal{P}_{{}_{\mathrm{B}}}^{0}$, i.e. we have a scale invariant amplitude for the power spectrum of the magnetic field in both the cases of $p=6$ and $4$. The second term in the expression within the square braces is the non-Gaussian contribution to the power spectrum of the magnetic field, which arises due to the interaction with the scalar perturbations. Note that the contribution is proportional to the product of the square of the non-Gaussianity parameter $b_{{}_{\mathrm{NL}}}({\bm{k}}_{1},{\bm{k}}_{2},{\bm{k}}_{3})$ and the scalar power spectrum $\mathcal{P}_{{}_{\mathrm{S}}}(k)$.
@@
×$\times\left(\frac{1}{y_{\mathrm{min}}}-\ln 2\right)\,.$ (92)
@@
It should be clear from the above expression that the relative difference is much larger than unity, mainly because of the presence of factor $1/(k\,\eta_{\mathrm{e}})^{4}$. Therefore, the non-Gaussian contribution to $\mathcal{P}_{{}_{\mathrm{B}}}(k)$ dominates over $\mathcal{P}_{{}_{\mathrm{B}}}^{0}$ in the case of $p=4$. Such a modified $\mathcal{P}_{{}_{\mathrm{B}}}^{{}_{\mathrm{M}}}(k)$ can be expressed as
@@ Appendix D Non-Gaussianity parameter in other cases
To further investigate the validity of the consistency relation (73), let us consider the ultra slow roll scenario wherein $p=6$ with $n=-1/3$. In this case, since $n\,p=-2$, the power spectrum of the magnetic field is not scale invariant but instead has a strong blue tilt and behaves as $k^{2}$ (i.e. $n_{{}_{\mathrm{B}}}=2$). Moreover, in this case, we find that the contributions to the cross-correlation from the first five terms ${\mathcal{B}}_{1}({\bm{k}}_{1},{\bm{k}}_{2},{\bm{k}}_{3})$ to ${\mathcal{B}}_{5}({\bm{k}}_{1},{\bm{k}}_{2},{\bm{k}}_{3})$ are of the same order. At the leading order, the resulting non-Gaussianity parameter can be obtained to be
@@
Figure 2: Density plots of the non-Gaussianity parameter $b_{{}_{\mathrm{NL}}}({\bm{k}}_{1},{\bm{k}}_{2},{\bm{k}}_{3})$ in the ultra slow roll scenario with $p=6$ and $n=-1/3$ (on the left) and slow roll inflation (on the right). In the ultra slow roll case, we have chosen the parameter values to be $\epsilon_{1\mathrm{i}}=10^{-3}$, $\eta_{\mathrm{i}}=-10^{6}\,\mathrm{Mpc}$ and $\eta_{\mathrm{e}}=-10^{-19}\,\mathrm{Mpc}$, as in the previous figure. To arrive at the results in the slow roll case, we have assumed the non-conformal coupling function to be of the form $J(\phi)$, as in the action (16). Moreover, we have assumed that $J\propto a^{2}$, corresponding to $\bar{n}=2$, which leads to a scale invariant power spectrum for the magnetic field. Also, we have chosen the values of the parameters to be $\epsilon_{1\mathrm{i}}=10^{-3}$, $H_{{}_{\mathrm{I}}}=1.25\times 10^{-5}\,M_{{}_{\mathrm{Pl}}}$ and $\eta_{\mathrm{e}}=-10^{-19}\,\mathrm{Mpc}$. Lastly, we have set $k_{3}$ to be the pivot scale $k_{\ast}$ in both the cases.
@@
Let us now compare the $b_{{}_{\mathrm{NL}}}({\bm{k}}_{1},{\bm{k}}_{2},{\bm{k}}_{3})$ we have obtained in the ultra slow roll scenarios of interest with the results arrived at in the standard slow roll inflationary scenario. In the latter case, as is often done, we shall assume that the electromagnetic field is described by the action (16) with a non-conformal coupling function of the form $J(\phi)$. For the slow roll case, in a background that is approximated to be that of de Sitter, and with a non-conformal coupling function that behaves as $J\propto a^{2}$ [i.e. when $\bar{n}=2$, cf. Eqs. (25) and (26)], we obtain a scale invariant power spectrum for the magnetic field (i.e. $n_{{}_{\mathrm{B}}}=0$). Also, in this case, the three-point cross-correlation of interest and the corresponding non-Gaussianity parameter can be easily calculated (for details, see Refs. [90, 101]). For suitable choices of parameters such that the inflationary scalar power spectrum is normalized to the value suggested by the CMB (and for $\eta_{\mathrm{e}}=-10^{-19}\,\mathrm{Mpc}$), we obtain the following estimates for $b_{{}_{\mathrm{NL}}}({\bm{k}}_{1},{\bm{k}}_{2},{\bm{k}}_{3})$ in the various limits of wave numbers (when evaluated at the pivot scale):
$b_{{}_{\mathrm{NL}}}^{\mathrm{eq}}(k_{\ast})$ $=180.18,$ (96a)
$b_{{}_{\mathrm{NL}}}^{\mathrm{fl}}(k_{\ast})$ $=2249.50,$ (96b)
@@
These numbers can be matched with the density plot in Fig. (2), where we have plotted $b_{{}_{\mathrm{NL}}}({\bm{k}}_{1},{\bm{k}}_{2},{\bm{k}}_{3})$ for an arbitrary triangular configuration of wave vectors. It is also clear from the above-mentioned value of $b_{{}_{\mathrm{NL}}}^{\mathrm{sq}}(k)$ that the consistency relation (73) is satisfied in this case (since $n_{{}_{\mathrm{B}}}=0$).
Finally, if we consider another slow roll case in an approximately de Sitter background with $J\propto a$ (corresponding to $\bar{n}=1$), we obtain that $\mathcal{P}_{{}_{\mathrm{B}}}(k)\propto k^{2}$ [i.e. $n_{{}_{\mathrm{B}}}=2$, cf. Eq. (32)]. Evaluating the three-point function in this scenario leads to the following values of the non-Gaussianity parameter $b_{{}_{\mathrm{NL}}}({\bm{k}}_{1},{\bm{k}}_{2},{\bm{k}}_{3})$ in the different limits of wave numbers:
$b_{{}_{\mathrm{NL}}}^{\mathrm{eq}}(k)$ $=\frac{23}{24}\simeq 0.96,$ (97a)
$b_{{}_{\mathrm{NL}}}^{\mathrm{fl}}(k)$ $=3,$ (97b)
@@
As in the preceding case, it is evident from the above value of $b_{{}_{\mathrm{NL}}}^{\mathrm{sq}}(k)$ that the consistency relation (73) is satisfied in this case as well.

### v2

@@ B.2 Case of $p=4$
$\quad-\,\left[9\,k_{1}^{4}-6\,k_{1}^{2}\,(k_{2}^{2}+k_{3}^{2})-3\,k_{2}^{4}-2\,k_{2}^{2}\,k_{3}^{2}-3\,k_{3}^{4}\right]\,\eta_{\mathrm{e}}^{4}\;\mathrm{Ei}[i\,(k_{1}+k_{2}+k_{3})\,\eta_{\mathrm{e}}]\biggr{)}.$ (84f)
We should mention that the expressions listed above have been arrived at without any approximations. In Sec. V, we have used these ${\mathcal{G}}_{C}({\bm{k}}_{1},{\bm{k}}_{2},{\bm{k}}_{3})$ to evaluate the contributions ${\mathcal{B}}_{C}({\bm{k}}_{1},{\bm{k}}_{2},{\bm{k}}_{3})$. We should add that, in listing the different contributions ${\mathcal{B}}_{C}({\bm{k}}_{1},{\bm{k}}_{2},{\bm{k}}_{3})$, we have retained only the leading order terms in $\eta_{\mathrm{e}}$.
@@ Appendix C Non-Gaussian contributions to $\mathcal{P}_{{}_{\mathrm{B}}}(k)$
We should mention here that similar analyses in the case of the contributions to the scalar power spectrum $\mathcal{P}_{{}_{\mathrm{S}}}(k)$ due to the scalar non-Gaussianity parameter $f_{{}_{\mathrm{NL}}}({\bm{k}}_{1},{\bm{k}}_{2},{\bm{k}}_{3})$ has been carried out earlier [117, 118, 50, 51, 119]. We shall suitably adopt the method followed in the case of the scalar power spectrum (in this context, see Refs. [51, 119]). Using the defining relation in Eq. (55), we can express the power spectrum of the magnetic field that is modified due to the non-Gaussianity parameter $b_{{}_{\mathrm{NL}}}({\bm{k}}_{1},{\bm{k}}_{2},{\bm{k}}_{3})$ to be
@@
In arriving at the second equation we have used the fact that $\mathcal{P}_{{}_{\mathrm{B}}}(k)=\mathcal{P}_{{}_{\mathrm{B}}}^{0}$, i.e. we have a scale-invariant amplitude for the power spectrum of the magnetic field in both the cases of $p=6$ and $4$. The second term in the expression within the square braces is the non-Gaussian contribution to the power spectrum of the magnetic field, which arises due to the interaction with the scalar perturbations. Note that the contribution is proportional to the product of the square of the non-Gaussianity parameter $b_{{}_{\mathrm{NL}}}({\bm{k}}_{1},{\bm{k}}_{2},{\bm{k}}_{3})$ and the scalar power spectrum $\mathcal{P}_{{}_{\mathrm{S}}}(k)$.
@@
×$\times\left(\frac{1}{y_{\mathrm{min}}}-\ln 2\right),$ (92)
@@
It should be clear from the above expression that the relative difference is much larger than unity, mainly because of the presence of the factor $1/(k\,\eta_{\mathrm{e}})^{4}$. Therefore, the non-Gaussian contribution to $\mathcal{P}_{{}_{\mathrm{B}}}(k)$ dominates over $\mathcal{P}_{{}_{\mathrm{B}}}^{0}$ in the case of $p=4$. Such a modified $\mathcal{P}_{{}_{\mathrm{B}}}^{{}_{\mathrm{M}}}(k)$ can be expressed as
@@ Appendix D Non-Gaussianity parameter in other cases
To further investigate the validity of the consistency relation given in Eq. (73), let us consider the ultra slow roll scenario wherein $p=6$ with $n=-1/3$. In this case, since $n\,p=-2$, the power spectrum of the magnetic field is not scale invariant but instead has a strong blue tilt and behaves as $k^{2}$ (i.e. $n_{{}_{\mathrm{B}}}=2$). Moreover, in this case, we find that the contributions to the cross-correlation from the terms ${\mathcal{B}}_{1}({\bm{k}}_{1},{\bm{k}}_{2},{\bm{k}}_{3})$, ${\mathcal{B}}_{2}({\bm{k}}_{1},{\bm{k}}_{2},{\bm{k}}_{3})$, ${\mathcal{B}}_{3}({\bm{k}}_{1},{\bm{k}}_{2},{\bm{k}}_{3})$, ${\mathcal{B}}_{5}({\bm{k}}_{1},{\bm{k}}_{2},{\bm{k}}_{3})$, and ${\mathcal{B}}_{6}({\bm{k}}_{1},{\bm{k}}_{2},{\bm{k}}_{3})$ are of the same order. At the leading order, the resulting non-Gaussianity parameter can be obtained to be
@@
Figure 2: Density plots of the non-Gaussianity parameter $b_{{}_{\mathrm{NL}}}({\bm{k}}_{1},{\bm{k}}_{2},{\bm{k}}_{3})$ in the ultra slow roll scenario with $p=6$ and $n=-1/3$ (on the left) and slow roll inflation (on the right). In the ultra slow roll case, we have chosen the parameter values to be $\epsilon_{1\mathrm{i}}=10^{-3}$, $\eta_{\mathrm{i}}=-10^{6}\,\mathrm{Mpc}$ and $\eta_{\mathrm{e}}=-10^{-19}\,\mathrm{Mpc}$, as in the previous figure. To arrive at the results in the slow roll case, we have assumed the non-conformal coupling function to be of the form $J(\phi)$, as in the action in Eq. (16). Moreover, we have assumed that $J\propto a^{2}$, corresponding to $\bar{n}=2$, which leads to a scale invariant power spectrum for the magnetic field. Also, we have chosen the values of the parameters to be $\epsilon_{1\mathrm{i}}=10^{-3}$, $H_{{}_{\mathrm{I}}}=1.25\times 10^{-5}\,M_{{}_{\mathrm{Pl}}}$ and $\eta_{\mathrm{e}}=-10^{-19}\,\mathrm{Mpc}$. Lastly, we have set $k_{3}$ to be the pivot scale $k_{\ast}$ in both the cases.
@@
Let us now compare the $b_{{}_{\mathrm{NL}}}({\bm{k}}_{1},{\bm{k}}_{2},{\bm{k}}_{3})$ we have obtained in the ultra slow roll scenarios of interest with the results arrived at in the standard slow roll inflationary scenario. In the latter case, as is often done, we shall assume that the electromagnetic field is described by the action in Eq. (16) with a non-conformal coupling function of the form $J(\phi)$. For the slow roll case, in a background that is approximated to be that of de Sitter, and with a non-conformal coupling function that behaves as $J\propto a^{2}$ [i.e. when $\bar{n}=2$, cf. Eqs. (25) and (26)], we obtain a scale invariant power spectrum for the magnetic field (i.e. $n_{{}_{\mathrm{B}}}=0$). Also, in this case, the three-point cross-correlation of interest and the corresponding non-Gaussianity parameter can be easily calculated (for details, see Refs. [92, 103]). For suitable choices of parameters such that the inflationary scalar power spectrum is normalized to the value suggested by the CMB (say, when $\epsilon_{1}\simeq 10^{-3}$ and $\eta_{\mathrm{e}}=-10^{-19}\,\mathrm{Mpc}$), we obtain the following estimates for $b_{{}_{\mathrm{NL}}}({\bm{k}}_{1},{\bm{k}}_{2},{\bm{k}}_{3})$ in the various limits of wave numbers (when evaluated at the pivot scale):
$b_{{}_{\mathrm{NL}}}^{\mathrm{eq}}(k_{\ast})$ $\simeq 180.18,$ (96a)
$b_{{}_{\mathrm{NL}}}^{\mathrm{fl}}(k_{\ast})$ $\simeq 2249.50,$ (96b)
@@
Note that in order to arrive at these numbers, we have neglected the contribution due to the term ${\mathcal{B}}_{4}({\bm{k}}_{1},{\bm{k}}_{2},{\bm{k}}_{3})$, since this term, being directly proportional to the first slow roll parameter $\epsilon_{1}$, would be subdominant compared to the other terms. These numbers can be matched with the density plot in Fig. 2, where we have plotted $b_{{}_{\mathrm{NL}}}({\bm{k}}_{1},{\bm{k}}_{2},{\bm{k}}_{3})$ for an arbitrary triangular configuration of wave vectors. It is also clear from the above-mentioned value of $b_{{}_{\mathrm{NL}}}^{\mathrm{sq}}(k)$ that the consistency relation in Eq. (73) is satisfied in this case (since $n_{{}_{\mathrm{B}}}=0$).
Finally, if we consider another slow roll case in an approximately de Sitter background with $J(\phi)\propto a$ (corresponding to $\bar{n}=1$), we obtain that $\mathcal{P}_{{}_{\mathrm{B}}}(k)\propto k^{2}$ [i.e. $n_{{}_{\mathrm{B}}}=2$, cf. Eq. (32)]. Evaluating the three-point function in this scenario leads to the following values of the non-Gaussianity parameter $b_{{}_{\mathrm{NL}}}({\bm{k}}_{1},{\bm{k}}_{2},{\bm{k}}_{3})$ in the different limits of wave numbers (when $\epsilon_{1}\simeq 10^{-3}$ and $\eta_{\mathrm{e}}=-10^{-19}\,\mathrm{Mpc}$):
$b_{{}_{\mathrm{NL}}}^{\mathrm{eq}}(k)$ $\simeq 0.96,$ (97a)
$b_{{}_{\mathrm{NL}}}^{\mathrm{fl}}(k)$ $\simeq 3,$ (97b)
@@
As in the preceding case, it is evident from the above value of $b_{{}_{\mathrm{NL}}}^{\mathrm{sq}}(k)$ that the consistency relation in Eq. (73) is satisfied in this case as well.